if 3^a-b = 1/9 and 3^a+b = 9 then a =
A) -2 b) 0 c) 1 d) 2 e) 3

Answer by stanbon(75887)

3^(a-b) = 1/9
3^(a+b) = 9
then a = ?
--------------------
3^(a-b) = 3^-2
3^(a+b) = 3^2
-------------------
a-b = -2
a+b = 2
---------
Adding you get:
2a = 0
a = 0
===================
Cheers,
Stan H.
Reply to stanbon@comcast.net


A) -2 b) 0 c) 1 d) 2 e) 3